Problem solving with equations
A rectangle has length 2cm longer than its width and its perimeter 52cm. Find its width

Answer by EdenWolf(517)

Let w = the width and w+2 = the length. Set up an equation.

2w+2(w+2)=52
2w+2w+4=52
4w+4=52
4w=48
w=12

The width is 12 cm.